OpenStudy (anonymous):

Simplify and write the result without negative exponents. (12a^8b^9/ 4a^7b^8)^-2

OpenStudy (anonymous):

You first have to bring every term to the -2 power. Then you move your terms to and from the numerator and denominator to eliminate the negatives. For example: x^4 is the same thing as 1/x^-4. Please respond if you need further clarification.

OpenStudy (anonymous):

1/(9 a^2 b^34)

OpenStudy (anonymous):

so i put 12 to the power of -2 making it into a decimal?

OpenStudy (anonymous):

is this \[(\frac{12a^8b^9}{4a^7b^8})^{-2}\]

OpenStudy (anonymous):

work inside the parentheses first to get \[(3ab)^{-2}\]

OpenStudy (anonymous):

yeah it is

OpenStudy (anonymous):

then the negative exponent means take the reciprocal, so write \[\frac{1}{(3ab)^2}\]

OpenStudy (anonymous):

if you want you can write \[\frac{1}{9a^2b^2}\]